OpenStudy (anonymous):

Carl deposited $1220 in a bank that pays 9% interest, compounded monthly. Find the amount he will have at the end of 3 years.

OpenStudy (anonymous):

FV = 1220 (1 + .09/12) ^ 12(3) FV = 1220 (1 + .0075)^ 36 FV = 1220 (1.0075 ^ 36) FV = 1220 (1.31) FV = 1598.20 I think I did that right. You might want to get a second opinion.

OpenStudy (anonymous):

it was right thank you

OpenStudy (anonymous):

I am glad....I would have been embarrassed if it had been wrong :)